How do you solve, in step by step form:
-(3/5)[2/3-1/2(3/4-1/2)]

Answer by ivy12003(22)

-(3/5)[2/3-1/2(3/4-1/2)]
=-(3/5)[2/3-1/2(3/4-2/4)]
=-(3/5)[2/3-1/2(1/4)]
=-(3/5)[2/3-1/8]
=-(3/5)[16/24-3/24]
=-(3/5)[13/24]
=-39/120 = -13/40